# Constrained Denoising, Empirical Bayes, and Optimal Transport

### v1

@@ 6.1 Astronomy
Our data comes from the Apache Point Observatory Galactic Evolution Experiment survey (APOGEE); see Majewski et al. (2017); Abolfathi et al. (2018); Ratcliffe et al. (2020); Soloff et al. (2024) for further detail. More specifically, we focus on Oxygen-Iron and Nitrogen-Iron relative abundances (denoted [O/Fe] and [N/Fe], respectively) for stars in the red clump catalog, following the analysis of Ratcliffe et al. (2020). For each star ($n=1{,}500$ stars sampled at random from the full catalog of $2.7\times 10^{4}$ stars), we have an estimate of the values [O/Fe] and [N/Fe], and the measurement error of these estimates is known. To state our model explicitly, we consider the following:
  • Every star $i$ has latent chemical abundances $\Theta_{i}=(\Theta_{i}^{\textnormal{[O/Fe]}},\Theta_{i}^{\textnormal{[N/Fe]}})$, and $\Theta_{1},\ldots,\Theta_{n}$ are i.i.d. from an unknown distribution $G$ on $\mathbb{R}^{2}$.
  • The measurement of every star $i$ admits error given by a diagonal covariance matrix $\Sigma_{i}$, the matrices $\Sigma_{1},\ldots,\Sigma_{n}$ are i.i.d. from some unknown distribution on $\mathbb{K}(2)$, and $\Theta_{1},\ldots,\Theta_{n}$ are independent of $\Sigma_{1},\ldots,\Sigma_{n}$.
  • For each star $i$, the observed pair of measurements $Z_{i}=(Z_{i}^{\textnormal{[O/Fe]}},Z_{i}^{\textnormal{[N/Fe]}})$ has conditional distribution $\mathcal{N}(\Theta_{i},\Sigma_{i})$ given $\Theta_{1},\ldots,\Theta_{n},\Sigma_{1},\ldots,\Sigma_{n}$, and $Z_{1},\ldots,Z_{n}$ are conditionally independent given $\Theta_{1},\ldots,\Theta_{n},\Sigma_{1},\ldots,\Sigma_{n}$.
The assumptions above are common in existing literature, e.g., each $\Sigma_{i}$ being diagonal (Ting and Weinberg, 2022, page 16), and $\Theta_{i}$ and $\Sigma_{i}$ being independent (Soloff et al., 2024, Section 1.1).
Next we describe the denoising methods of interest. All of our approaches will be based on smooth $G$-modeling, as in Example 4.2, where we assume that $G$ is a Gaussian location mixture model, where the covariance matrix of each Gaussian component is bounded below, in the PSD order, by
$\begin{pmatrix}\tau^{2}_{\textnormal{[O/Fe]}}&0\\ 0&\tau^{2}_{\textnormal{[N/Fe]}}\\ \end{pmatrix},\qquad\textnormal{ where }\quad\tau^{2}_{\textnormal{[O/Fe]}}=2.0699\times 10^{-3}\textnormal{ and }\tau^{2}_{\textnormal{[N/Fe]}}=8.1521\times 10^{-4}.$
The diagonal elements above were chosen as the 20% empirical quantile of the diagonal elements in the observed heteroskedasticity, $\Sigma^{\textnormal{[O/Fe]}}_{1},\ldots\Sigma^{\textnormal{[O/Fe]}}_{n}$. Then, we compute the heteroskedastic multivariate NPMLE of Soloff et al. (2024) over this restricted class of $G$, yielding an unconstrained EB denoiser $\hat{\delta}_{\mathcal{B}}$. Last, we use $\hat{\delta}_{\mathcal{B}}$ to compute the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and the marginal distribution-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{DCB}}$, as described in Section 5.
@@
The results are shown in Figure 6.1, which deserves some discussion. First, note that the EB denoiser $\hat{\delta}_{\mathcal{B}}$ recovers some interesting latent low-dimensional structure from the data. However, the difference between the EB denoiser $\hat{\delta}_{\mathcal{B}}$ and the variance-constrained EB denoiser $\hat{\delta}_{\mathcal{VCB}}$ illustrates that the shrinkage of $\hat{\delta}_{\mathcal{B}}$ may yield overestimation of this latent structure. Also, we note that the transformation from the EB denoiser $\hat{\delta}_{\mathcal{B}}$ to the variance-constrained EB denoiser $\hat{\delta}_{\mathcal{VCB}}$ is an affine but it is not merely scaling; the upper branch slightly changes its angle, which suggests that the correlation between the two components may be underestimated by the EB denoiser. Lastly, we note that the distribution-constrained EB denoiser $\hat{\delta}_{\mathcal{DCB}}$ reveals a similar latent low-dimensional structure, but that the resulting distribution is more equispaced and this allows one identify more subtle features in the distribution (e.g., the shape of the cluster at the elbow, the varying density profile along the two branches, and the separated clusters appearing at the ends of the branches).
@@ 6.2 Baseball
Our data consists of all batters in their first season of major league play during the 2022 to 2024 season of the MLB which is publicly available through the website FanGraphs. [^2] Available at: https://www.fangraphs.com/ For each player ($n=324$), we observe his RBIs and his number of games played, in both their final minor league season and their initial major league season. To state our model explicitly, we consider the following:
  • Every player $i$ has a latent bivariate skill $\Theta_{i}=(\Theta_{i}^{\textnormal{min}},\Theta_{i}^{\textnormal{maj}})$, and $\Theta_{1},\ldots,\Theta_{n}$ are i.i.d. from an unknown distribution $G$ on $[0,\infty)\times[0,\infty)$.
  • Every player $i$ participates in a random proportion $\lambda_{i}=(\lambda_{i}^{\textnormal{min}},\lambda_{i}^{\textnormal{maj}})$ of all games per minor- and major-league season, and $\lambda_{1},\ldots,\lambda_{n}$ are i.i.d. from an unknown distribution on $[0,\infty)\times[0,\infty)$, and $\lambda_{1},\ldots,\lambda_{n}$ are independent of $\Theta_{1},\ldots,\Theta_{n}$.333 The assumption of independence between $\Theta_{1},\ldots,\Theta_{n}$ and $\lambda_{1},\ldots,\lambda_{n}$ is debatable and warrants further investigation in future work. On the one hand, good batters are more likely to be placed in a batting lineup than bad batters. On the other hand, players who are desired for non-batting skills (catching, fielding, etc.) must be placed in the batting lineup in order to exercise these other skills.
  • For each plater $i$, his RBI count $Z_{i}=(Z_{i}^{\textnormal{min}},Z_{i}^{\textnormal{maj}})$ has conditional distribution $\textnormal{Poi}(\lambda_{i}^{\textnormal{min}}\Theta_{i}^{\textnormal{min}})\otimes\textnormal{Poi}(\lambda_{i}^{\textnormal{maj}}\Theta_{i}^{\textnormal{maj}})$ given $\Theta_{1},\ldots,\Theta_{n},\lambda_{1},\ldots,\lambda_{n}$, and $Z_{1},\ldots,Z_{n}$ are conditionally independent given $\Theta_{1},\ldots,\Theta_{n},\lambda_{1},\ldots,\lambda_{n}$.
Next we describe our denoising methods. All of our methods are based on $G$-modeling (without smoothness assumptions) where we posit that $G$ a bivariate heterogeneous Poisson mixture model with conditionally independent components; this is similar to existing studies on the univariate Poisson case (Polyanskiy and Wu, 2021; Jana et al., 2024) and the multivariate heteroskedastic Gaussian case (Soloff et al., 2024). First, we compute the NPMLE for $G$, yielding an unconstrained EB denoiser $\hat{\delta}_{\mathcal{B}}$. Second, as outlined in Section 5, we compute the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and the marginal general-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{GCB}}$ corresponding to constraints on the first and second moment and on component-wise nonnegativity, as in Example 3.11.
@@ Acknowledgments
We would like to thank Stephen Raudenbush and Tom Loredo for helpful discussions and for pointing us to the literature on constrained EB estimation. We also thank Ellen Woods for many useful conversations about baseball.
Appendix A Computational Considerations
In this section, we briefly describe some computational considerations that arise in the implementation of various algorithms described in the main body of the paper. Our procedures also inherit some computational complexity from the unconstrained EB denoisers which they take as input. Convex optimization problems are solved via CVXPY [Diamond and Boyd, 2016] using the Mosek  [ApS, 2025] interior point convex programming solver.
A.1 $G$-modeling via nonparametric maximum likelihood
  1. 1.
  2. 2.
In the Gaussian models (Figure 1.1, Figure 5.1, and Figure 6.1) we inialize the NPMLE atoms in Step 1 to be the data points themselves, following the so-called exemplar$+$ method of Soloff et al. [2024]. In the Poisson model of Figure 6.2 we initialize the NPMLE atoms of step 1 to be a $50\times 50$ equispaced grid for the minimal bounding box of the data.
A.2 Variance-Constrained Denoising
Second, we consider the two variance-constrained denoising problems in the heterogeneous setting of Section 5. The marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ does not require any further computational burden compared to $\hat{\delta}_{\mathcal{VCB}}$, at least in the case where $\hat{\delta}_{\mathcal{B}}$ arises from $G$-modeling. This is the case in all of our applications except for Figure 5.1. The conditional variance-constrained denoiser $\hat{\delta}_{\mathcal{C}\mathcal{VCB}}$ is more complicated since it requires the numerical integration of the terms in (5.6) for each evaluation of $\hat{\delta}_{\mathcal{C}\mathcal{VCB}}(Z_{i},\Xi_{i})$; we compute these integrals via Monte Carlo with 100 trials each.
A.3 Distribution- and General-Constrained Denoising
Our procedures Algorithm 2 and Algorithm 3 for approximating the solution to the distribution-constrained problem ($\mathcal{DCB}$) and the general-constrained problem ($\mathcal{GCB}$) both require numerically solving a linear program (lines 7 to 10), [^5] Techniques from computational optimal transport Peyré and Cuturi [2019], e.g., the Sinkhorn algorithm, would make this step more efficient. We did not pursue this here. so we briefly discuss some details of this. First note that some discretization of the $\eta\in\mathbb{R}^{m}$ coordinate is necessary, and we consequently solve this problem in CVXPY using the solver Mosek.
We always choose the discretization to be an equispaced grid on some axis-aligned bounding box, and the main difficulty is choosing an appropriate discretization.
In the cases where $m=1$ (i.e., Figure 5.1), the complexity of the linear program problem scales linearly in the number of discretization points $k$. So, it is possible to take $k$ quite large without computational burden; for the sake of simplicity. In these cases, we fixed the discretization of $\eta$ to consist of $k=200$ equispaced points in the minimal bounding interval of the observations.
In the cases where $m\geq 2$ (i.e., Figure 1.1, Figure 6.2, and Figure 6.1) we need to resort to some heuristics since the complexity of the resulting linear program scales poorly with the discretization. First, we take a discretization of $\eta$ to consist of $k^{m}$ many equispaced points in the minimal bounding box of the observations. Then, we sample (without replacement) $K$ many gridpoints according to probability they are assigned under $\hat{G}_{n}$. This is like a quasi-Monte Carlo estimate of $\hat{G}_{n}$ since subsampling a grid provides increased regualrity over direct sampling from $\hat{G}_{n}$. In these cases, we took $k=200$ and $K=1000$.
Additionally, we remark that naively discretizing the linear program may lead to infeasibility; this is because $\hat{G}_{n}$ witnesses feasibility for arbitrary probability measures but not necessarily for probability measures supported on the prescribed grid points. To avoid this, we augment the discretization by adding the NPMLE atoms in addition to the grid points.
Appendix B Additional Formulas for Example 4.4
Previously we saw that, if $\{P_{\theta}\}_{\theta}$ is a suitable univariate exponential family and $G$ is a conjugate prior, then the Bayes denoiser $\delta_{\mathcal{B}}$ is a linear function whose slope and intercept an be estimated from the marginal distribution of $Z$. Consequently, the EB approximation $\hat{\delta}_{\mathcal{B}}$ of the Bayes denoiser $\delta_{\mathcal{B}}$ admits a closed-form solution. Similarly, the EB approximation $\hat{\delta}_{\mathcal{VCB}}$ of the variance-constrained Bayes denoiser $\delta_{\mathcal{VCB}}$ also admits a closed-form solution. In this section, we give some explicit formulas for $\hat{\delta}_{\mathcal{B}}$ and $\hat{\delta}_{\mathcal{VCB}}$ for some parametric models of interest.
Recall that $\{P_{\theta}\}_{\theta}$ is assumed to be a mean-parameterized exponential family with a quadratic variance function; that is, there exists a quadratic polynomial $V:\mathbb{R}\to\mathbb{R}$ such that we have
$\mathbb{E}[Z\,|\,\Theta]=\Theta\qquad\textnormal{and}\qquad\textnormal{Var}(Z\,|\,\Theta)=V(\Theta)$
almost surely.
In this case, the Bayes denoiser is
@@ A.3 Distribution- and General-Constrained Denoising
where $\mu=\mathbb{E}[\Theta]$ and $a^{2}=\textnormal{Var}(\Theta)$ are the mean and variance of $G$. Now note that the law of total variance shows that the marginal variance $s^{2}:=\textnormal{Var}(Z)$ satisfies the identity $s^{2}=(1+\frac{1}{2}V^{\prime\prime}(0))a^{2}+V(\mu)$, hence an alternative expression for $\delta_{\mathcal{B}}$ is
$\delta_{\mathcal{B}}(z)=\frac{1-s^{-2}V(\mu)}{1+\frac{1}{2}V^{\prime\prime}(0)}z+\left(1-\frac{1-s^{-2}V(\mu)}{1+\frac{1}{2}V^{\prime\prime}(0)}\right)\mu,$
which depends only on $\mu$ and $s^{2}$. As such, a natural approximation $\delta_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$ is the plug-in estimator
@@
where $\hat{\mu}:=\frac{1}{n}\sum_{i=1}Z_{i}$ and $\hat{\sigma}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(Z_{i}-\hat{\mu})^{2}$ are the empirical mean and variance of $Z_{1},\ldots,Z_{n}$.
By combining Theorem 3.3 with the identity $s^{2}=(1+\frac{1}{2}V^{\prime\prime}(0))a^{2}+V(\mu)$, it is easy to see that the variance-constrained Bayes denoiser is given by
@@
hence a natural EB approximation for it is given by
$\hat{\delta}_{\mathcal{VCB}}(z)=\sqrt{\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}}z+\left(1-\sqrt{\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}}\right)\hat{\mu},$
by the plug-in principle.
In the remainder of this section, we give explicit formulas for the denoisers above, for the usual parametric models of interest.
@@
Appendix C Discussion of Rates of Convergence
C.1 Variance-Constrained Denoising

### v2

@@ 6.1 Astronomy
Our data comes from the Apache Point Observatory Galactic Evolution Experiment survey (APOGEE); see Majewski et al. (2017); Abolfathi et al. (2018); Ratcliffe et al. (2020); Soloff et al. (2024) for further detail. More specifically, we focus on Oxygen-Iron and Nitrogen-Iron relative abundances (denoted [O/Fe] and [N/Fe], respectively) for stars in the red clump catalog, following the analysis of Ratcliffe et al. (2020). For each star ($n=2{,}000$ stars sampled at random from the full catalog of $2.7\times 10^{4}$ stars), we have an estimate of the values [O/Fe] and [N/Fe], and the measurement error of these estimates is known.
To state our model explicitly, we make the following modeling assumptions: 1. Every star $i$ has latent chemical abundances $\Theta_{i}=(\Theta_{i}^{\textnormal{[O/Fe]}},\Theta_{i}^{\textnormal{[N/Fe]}})$, and $\Theta_{1},\ldots,\Theta_{n}$ are i.i.d. from an unknown distribution $G$ on $\mathbb{R}^{2}$. 2. The measurement of every star $i$ admits error given by a diagonal covariance matrix $\Sigma_{i}$, the matrices $\Sigma_{1},\ldots,\Sigma_{n}$ are i.i.d. from some unknown distribution on $\mathbb{K}(2)$, and $\Theta_{1},\ldots,\Theta_{n}$ are independent of $\Sigma_{1},\ldots,\Sigma_{n}$. 3. For each star $i$, the observed pair of measurements $Z_{i}=(Z_{i}^{\textnormal{[O/Fe]}},Z_{i}^{\textnormal{[N/Fe]}})$ has conditional distribution $\mathcal{N}(\Theta_{i},\Sigma_{i})$ given $\Theta_{1},\ldots,\Theta_{n},\Sigma_{1},\ldots,\Sigma_{n}$, and $Z_{1},\ldots,Z_{n}$ are conditionally independent given $\Theta_{1},\ldots,\Theta_{n},\Sigma_{1},\ldots,\Sigma_{n}$. The assumptions above are common in existing literature, e.g., each $\Sigma_{i}$ being diagonal (Ting and Weinberg, 2022, page 16), and $\Theta_{i}$ and $\Sigma_{i}$ being independent (Soloff et al., 2024, Section 1.1).
Next we describe the denoising methods of interest. All of our approaches will be based on smooth $G$-modeling, as in Example 4.2, where we assume that $G$ is a Gaussian location mixture model, where the covariance matrix of each Gaussian component is bounded below, in the PSD order, by $\tau^{2}I_{2}$, where $\tau^{2}=0.0025$. Then, we compute the heteroskedastic multivariate NPMLE of Soloff et al. (2024) over this restricted class of $G$, yielding an unconstrained EB denoiser $\hat{\delta}_{\mathcal{B}}$. Last, we use $\hat{\delta}_{\mathcal{B}}$ to compute the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and the marginal distribution-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{DCB}}$, as described in Section D.
@@
The results are shown in Figure 6.1, which deserves some discussion. First, note that the EB denoiser $\hat{\delta}_{\mathcal{B}}$ recovers some interesting latent low-dimensional structure from the data. However, the difference between the EB denoiser $\hat{\delta}_{\mathcal{B}}$ and the variance-constrained EB denoiser $\hat{\delta}_{\mathcal{VCB}}$ illustrates that the shrinkage of $\hat{\delta}_{\mathcal{B}}$ may yield overestimation of this latent structure. Also, we note that the transformation from the EB denoiser $\hat{\delta}_{\mathcal{B}}$ to the variance-constrained EB denoiser $\hat{\delta}_{\mathcal{VCB}}$ is an affine but it is not merely scaling; the upper branch slightly changes its angle, which suggests that the correlation between the two components may be underestimated by the EB denoiser. Lastly, we note that the distribution-constrained EB denoiser $\hat{\delta}_{\mathcal{DCB}}$ reveals a similar latent low-dimensional structure, but that the resulting distribution is more equispaced and this allows one identify more subtle features in the distribution (e.g., the separated cluster appearing at the ends of the top branch, and the positions of points outside the bulk of the data set).
@@ 6.2 Baseball
Our data consists of all batters in their first season of major league play during the 2022 to 2024 season of the MLB which is publicly available through the website FanGraphs. [^3] Available at: https://www.fangraphs.com/ For each player ($n=324$), we observe his RBIs and his number of games played, in both their final minor league season and their initial major league season. To state our model explicitly, we make the following modeling assumptions: 1. Every player $i$ has a latent bivariate skill $\Theta_{i}=(\Theta_{i}^{\textnormal{min}},\Theta_{i}^{\textnormal{maj}})$, and $\Theta_{1},\ldots,\Theta_{n}$ are i.i.d. from an unknown distribution $G$ on $[0,\infty)\times[0,\infty)$. 2. Every player $i$ participates in a random proportion $\lambda_{i}=(\lambda_{i}^{\textnormal{min}},\lambda_{i}^{\textnormal{maj}})$ of all games per minor- and major-league season, and $\lambda_{1},\ldots,\lambda_{n}$ are i.i.d. from an unknown distribution on $[0,\infty)\times[0,\infty)$, and $\lambda_{1},\ldots,\lambda_{n}$ are independent of $\Theta_{1},\ldots,\Theta_{n}$.444 The assumption of independence between $\Theta_{1},\ldots,\Theta_{n}$ and $\lambda_{1},\ldots,\lambda_{n}$ is debatable and warrants further investigation in future work. On the one hand, good batters are more likely to be placed in a batting lineup than bad batters. On the other hand, players who are desired for non-batting skills (catching, fielding, etc.) must be placed in the batting lineup in order to exercise these other skills. 3. For each plater $i$, his RBI count $Z_{i}=(Z_{i}^{\textnormal{min}},Z_{i}^{\textnormal{maj}})$ has conditional distribution $\textnormal{Poi}(\lambda_{i}^{\textnormal{min}}\Theta_{i}^{\textnormal{min}})\otimes\textnormal{Poi}(\lambda_{i}^{\textnormal{maj}}\Theta_{i}^{\textnormal{maj}})$ given $\Theta_{1},\ldots,\Theta_{n},\lambda_{1},\ldots,\lambda_{n}$, and $Z_{1},\ldots,Z_{n}$ are conditionally independent given $\Theta_{1},\ldots,\Theta_{n},\lambda_{1},\ldots,\lambda_{n}$.
Next we describe our denoising methods. All of our methods are based on $G$-modeling (without smoothness assumptions) where we posit that $G$ a bivariate heterogeneous Poisson mixture model with conditionally independent components; this is similar to existing studies on the univariate Poisson case (Polyanskiy and Wu, 2021; Jana et al., 2024) and the multivariate heteroskedastic Gaussian case (Soloff et al., 2024). First, we compute the NPMLE for $G$, yielding an unconstrained EB denoiser $\hat{\delta}_{\mathcal{B}}$. Second, as outlined in Section D, we compute the marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ and the marginal general-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{GCB}}$ corresponding to constraints on the first and second moment and on component-wise nonnegativity.
@@
7 Discussion
Constrained denoising is by now a classical idea that has found use across applications. The optimal transport framework developed in this paper substantially extends the scope of constrained denoising by showing that it is fully compatible with nonparametric empirical Bayes estimators, extending it to higher-dimensional settings, and introducing constraints beyond variance—including distributional and general constraints, as well as conditional constraints under heterogeneity.
For practitioners, several natural questions arise: Which empirical Bayes denoiser should be used among a suite of parametric and nonparametric options? Which type of constraint is most appropriate—variance, distributional, or general? And if general constraints are chosen, which functionals of the latent distribution should be matched? We do not provide definitive answers to these questions, as the optimal choices depend on both the specific application domain and the goals of downstream analyses. Lee et al. (2024) conducted a detailed study providing guidelines for choosing among denoising methods (including variance-constrained approaches) for the analysis of multisite trials. With the expanded framework developed here, we envisage similar guidelines emerging for other concrete application areas, guided by the requirements of domain experts.
From the point of view of statistical theory, an important outstanding question (which is beyond the scope of this paper) concerns lower bounds for our rates of convergence in Section 4. That is, are the rates of convergence in our Theorem 4.3 and Theorem 4.4 optimal? We believe such questions will be interesting for future work, especially in the setting of a nonparametric latent variable distribution and a Gaussian likelihood.
Reproducibility
All figures in the paper can be reproduced in Python with the Jupyter notebooks available at https://github.com/aqjaffe/constrained-denoising-EB-OT.
@@ Acknowledgments
We would like to thank Stephen Raudenbush and Tom Loredo for helpful discussions and for pointing us to the literature on constrained EB estimation. We also thank Brad Ross for providing detailed comments on an early version of this manuscript. Finally, we thank Ellen Woods for many useful conversations about baseball.
Appendix A Additional Empirical Bayes Approaches
While the main body of the paper is primarily focused on nonparametric EB methods, we note that our results apply to (unconstrained) EB denoisers $\hat{\delta}_{\mathcal{B}}$ arising from many different approaches. In this section we briefly review some alternative approaches which we will believe will be important in other applications.
A.1 $F$-Modeling
In some settings, the Bayes denoiser $\delta_{\mathcal{B}}$ depends only on the marginal distribution $F$ of $Z_{1},\ldots,Z_{n}$. Then an EB approximation $\hat{\delta}_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$ arises by modeling $F$ directly, rather than $G$. This is possible when $P$ is a continuous exponential family with natural parameterization, due to Eddington’s/Tweedie’s formula [Dyson, 1926, Efron, 2011], but it is also possible in many special cases where similar formulas are available [Robbins, 1956, 1982, Cressie, 1982]. Here we follow Efron [2014] and distinguish between $G$-modeling and $F$-modeling approaches, although we note that $F$-modeling can also be viewed as a special case of $G$-modeling.
A.2 Conjugate Parametric Models
Suppose in $d=m=1$ that $\{P_{\theta}\}_{\theta}$ is a mean-parameterized exponential family and that $G$ is a conjugate prior for $\{P_{\theta}\}_{\theta}$. We also assume that $\{P_{\theta}\}_{\theta}$ has a quadratic variance function, meaning there exists a quadratic polynomial $V:\mathbb{R}\to\mathbb{R}$ such that $P_{\theta}$ has variance $V(\theta)$ for all $\theta$; see Morris [1982] for more on such models. It is well-known (see Diaconis and Ylvisaker [1979]) that the Bayes denoiser $\delta_{\mathcal{B}}$ is a linear function in this setting; we have
@@ A.2 Conjugate Parametric Models
where $\hat{\mu}:=\frac{1}{n}\sum_{i=1}Z_{i}$ and $\hat{s}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(Z_{i}-\hat{\mu})^{2}$ are the empirical mean and variance of $Z_{1},\ldots,Z_{n}$. One can easily show in this setting that the convergence of $\hat{\delta}_{\mathcal{B}}$ to $\delta_{\mathcal{B}}$ has the parametric rate $n^{-1}$.
In this setting, we may use Theorem 3.3 to see that the variance-constrained Bayes denoiser is given explicitly by
@@
and also that $\hat{\delta}_{\mathcal{VCB}}$ is given by
$\hat{\delta}_{\mathcal{VCB}}(z)=\sqrt{\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}}z+\left(1-\sqrt{\frac{1-\hat{s}^{-2}V(\hat{\mu})}{1+\frac{1}{2}V^{\prime\prime}(0)}}\right)\hat{\mu}.$
In the remainder of this section, we give some explicit formulas for the denoisers above, for the usual parametric models of interest.
@@
Appendix B Discussion of Rates of Convergence
B.1 Variance-Constrained Denoising